if 3x-2[3x+1/2] =1 then x=?
Answer by stanbon(75887)

3x-2[3x+1/2] = 1
-------------
3x -6x - 1 = 1
-----
-3x = 2
x = -2/3
----------------
Cheers,
Stan H.
================